OpenStudy (anonymous):

6. Which situation could be solved using this expression? 2k – 3 A. Tamara made three times as many free throws as Marla. Marla made 3 free throws. How many free throws did Tamara make? B. Tamara made 3 fewer than twice the number of free throws Marla made. How many free throws did Tamara make? C. Tamara made 3 fewer than twice the number of free throws Marla made. How many free throws did Marla make? D. Tamara made 2 free throws. Marla made 3 free throws. Jenny made some free throws. How many free throws did they make altogether?

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

anyone help me on this question please?

OpenStudy (jdoe0001):

2k -3 2k if Marla made "k" free throws, that means twice as much will be "2k" less 3 from that will be 2k - 3 so Tamara made twice as much, less 3 of that or 3 less than twice as much as Marla

OpenStudy (anonymous):

is it A,B,C,D?

OpenStudy (jdoe0001):

well, B and C are about the same, the names just swap about

OpenStudy (anonymous):

wow you are really good at math

OpenStudy (jdoe0001):

t = Tamara m =Marla let's say see option B t = 2k -3 How many free throws did Tamara make? well, she made 2k -3

OpenStudy (anonymous):

Evaluate. 6 • [7 + (15 – 8)] ÷ 3

OpenStudy (jdoe0001):

well, that's just a PEMDAS matter

OpenStudy (jdoe0001):

grouped elements are done first, from innermost, to outer

OpenStudy (jdoe0001):

then multiplicaton and division from left to right

OpenStudy (anonymous):

can you tell me the answer

OpenStudy (anonymous):

is it 2.33333333333?

OpenStudy (anonymous):

or 16.33333333333

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

\(\Large \bf \color{red}{6 \cdot [\color{blue}{7 + (\color{green}{15 - 8})}] \div 3}\) you'd do the "green" part first the "blue" 2nd the "red" or outermost last

OpenStudy (anonymous):

ok thank you so much

OpenStudy (jdoe0001):

15-8 = 7 [ 7 + (7)] = 14 \(\bf 6 \times 14 = 84\\ 84 \div 3 = 28\)

OpenStudy (anonymous):

Translate the phrase into a math expression. twice a number decreased by 12 A. 2 · (n – 12) B. 2 · n - 12 C. 2 · n ÷ 12 D. 12 – 2 · n